# MOA-2010-BLG-328: Keck and HST Expose the Limits of Occams Razor in Microlensing

### v1

@@ 4 Image Constrained Light Curve Models
In the following sub-sections we discuss each higher order effect that is included in the light-curve modeling of this event. The best fit parameters for these effects are shown in Table 4, with the values and uncertainties for the Einstein radius crossing time ($t_{\rm E}$),time of closest approach ($t_{0}$), source crossing time ($t_{*}$), impact parameter ($u_{0}$), the instantaneous projected separation which is scaled to $\theta_{\rm E}$ ($s$), the host-planet mass ratio ($q$), and the Einstein ring radius ($\theta_{\rm E}$).
@@ 4.0.2 Orbital Motion of Lens System
Orbital motion of the lens system is always present, however, their positions are often assumed to be fixed during the microlensing event. This is because a microlensing event’s duration is often on the order of a few days to weeks, making is a small fraction of the lens systems’ orbital period of several years. In our model we assume a face-on circular orbit with the parameters of orbital angular speed, $\omega$, and the separation rate of $ds/dt$. If microlensing parallax effects are present on the light-curve, then it is physically realistic that the model include orbital motion effects too. However, due to rotational symmetry, $\pi_{\rm E,E}$ (which is the component perpendicular to the instantaneous direction of the Earth’s acceleration) is often degenerate with $\omega$ (Batista et al., 2011; Skowron et al., 2011).
@@ 4.0.3 Binary Source
The parameters due to the binary source effect are the time of closest approach ($t_{0,2}$) and impact parameter ($u_{0,2}$) of the second source. Additionally, we have the $I$ and $V$ band flux ratios between the second source ($S_{2}$) and the primary source ($S_{1}$), fr2I and fr2V, respectively. The xallarap effect is when the source orbital motion due to the second source invokes the parameters $\rm dt_{\rm E21}$, which is implied by $t_{\rm E2}-t_{\rm E1}$; and $d\theta$, implied by $\theta_{\rm E2}-\theta_{\rm E1}$.
The light curve is plotted in Figure 5. It does not include the SAAO and $\mu$Fun V-band data due to the two sources not being the same color. From our best fit model we find that the primary and secondary source magnitude and color to be $(V-I,I)_{S}=(2.189,19.609)\pm(0.041,0.023)$ and $(V-I,I)_{S2}=(1.918,20.864)\pm(0.373,0.131)$, respectively. These are found using the constraints on the total magnitude and color from the 2018 HST data.
Figure 5: Best fit 2L2S light curve for MB10328 using constraints from the Keck and HST high resolution follow-up data. The parameters are described in Section 4 and Table 4.
5 Lens System Parameters
@@
The physical properties of the lens system can be found through a series of mass-distance relations which include the angular Einstein ring radius, $\theta_{\rm E}$, or the microlensing parallax $\pi_{\rm E}$ (Beaulieu et al., 2008; Gaudi, 2012). These relations are:
@@
where $\pi_{\rm rel}=\rm AU(\frac{1}{D_{\rm L}}-\frac{1}{D_{\rm S}}),$ and $\kappa=4G/(c^{2}AU)=8.144\ \rm mas\ M_{\odot}^{-1}$. Therefore, with Equations 3 and 4 an expression can be found for the lens mass $M_{\rm L}$, with no dependance on the lens or source distance, $D_{\rm L}$ or $D_{\rm S}$, respectively. The lens distance, however, does depend on the source distance.
For the event MB10328, a direct measurement of the lens-source proper motion has been made, therefore $\theta_{E}$ can be constrained with $\theta_{E}=\mu_{\rm rel,G}t_{E}$. However, $\mu_{\rm rel,G}$ is in the instataneous geocentric inertial reference frame which moves with the earth at peak magnification. Since the proper motion derived from the high angualr resolution follow up images are in the heliocentric frame, it needs to be converted to the geocentric one. This can be done using the equation:
${\boldmath\mu}_{\rm rel,G}=\mu_{\rm rel,Hel}-\frac{v_{\oplus}\pi_{\rm rel}}{\rm AU}$ (5)
Where $v_{\oplus}$ is the projected velocity of the earth relative to the sun at the time of the events peak magnification. For MB10328, this was $v_{\oplus E,N}=(28.878,-0.689)$ km/sec $=(6.088,-0.145)$ AU/yr at HJD’ = 5406. Since $\pi_{\rm rel}$ is the relative parallax, we can rewrite equation 5 as:
${\boldmath\mu}_{\rm rel,G}=\mu_{\rm rel,Hel}-(6.088,-0.145)\times(1/D_{\rm L}-1/D_{\rm S}),$ (6)
where $\mu_{\rm Hel}$ and $\mu_{\rm rel,G}$ are in units of mas/yr and $D_{\rm L}$ and $D_{\rm S}$ are in units of kpc. This relation can be used in our Bayesian analysis of the light curve (Section 4) alongside our and high resolution follow-up constraints from Keck and $HST$. Therefore, we determine the geocentric relative proper motion to be $\mu_{\rm rel,G}=3.728\pm 0.145\ \rm mas\ yr^{-1}$.
@@
$m_{\rm L}=-5+5\log_{10}(D_{\rm L}/1\ \rm kpc)+A_{K,L}+M_{\mathcal{K}}(M_{L}),$ (7)
where $M_{\mathcal{K}}(M_{\rm L})$ is a $K$ band luminosity relation. The interstellar K band extinction along the lens line of sight is given by $A_{\rm K,L}$. For this event the extinction we use is $A_{\rm K,L}=0.185$ based on the Surot et al. (2020) E(J-K) map and Nishiyama et al. (2009) IR extinction law. The empirical mass-luminosity relations are from Henry & McCarthy (1993) ($M_{\rm L}\geqslant 0.66\ M_{\odot}$), Delfosse et al. (2000) ($0.54\ M_{\odot}\geqslant M_{\rm L}\geqslant 0.12\ M_{\odot}$) and Henry et al. (1999) ($0.10\ M_{\odot}\geqslant M_{\rm L}\geqslant 0.07\ M_{\odot}$). They are also used in Bennett et al. (2015, 2016, 2018). The empirical mass-luminosity relation merges these three relations, smoothly interpolating between them.
By summing over the remodeled light curve MCMC results that are presented in Table 4, we can determine the posterior distribution of the lens system properties. A common assuption used in the Bayesian analysis of microlensing events is that all stars are equally likely to host a planet with a fixed mass ratio. However, more recent studies show that planet hosting probability of a fixed mass ratio increases linearly with the host mass. Therefore, this prior is included when summing over the MCMC results. Since we have constraints from high resolution imaging, this prior has a small effect on the results.
From the Bayesian probability distribution we find that the host star is in the disk at $D_{\rm L}=3.828\pm 1.113$ kpc with a mass of $M_{\rm L}=0.565\pm 0.184\ M_{\odot}$ and a planetary companion with $M_{\rm p}=136.9\pm 72.76\ M_{\rm Earth}$. Using $\alpha_{\perp}=\rm sD_{L}\theta_{E}$, where $s$ is the projected separation from the light curve, we find the planet’s 2D separation to be $\alpha_{\perp}=3.374\pm 0.770$ AU. However, the distribution is not a single Gaussian, and when looking at the probability distributions in Figure 6 we see a clear degeneracy in these physical parameters, which we investigate in more detail in the following section 5.1.
Table 5: Planetary system parameters from high angular resolution follow up constraints compared to those found by Furusawa et al. (2013).
Parallax plus orbital motion This Work
Parameter ($u_{0}<0$) ($u_{0}>0$) Xallarap $\pi_{\rm E}<\pi_{\rm E,N}-\pi_{\rm E,E}-0.33$ $\pi_{\rm E}>\pi_{\rm E,N}-\pi_{\rm E,E}-0.33$
q/$10^{-4}$ $2.60\pm 0.53$ $3.68\pm 1.26$ $5.16\pm 0.06$ 7.608 $\pm$ 1.621 3.156 $\pm$ 0.232
$\theta_{\rm E}$ (mas) 0.98 $\pm$ 0.12 0.83 $\pm$ 0.14 0.68 $\pm$ 0.04 0.701 $\pm$ 0.020 0.899 $\pm$ 0.322
$\mu_{\rm rel,G}$ ($\rm mas\ yr^{-1}$) 5.71 $\pm$ 0.70 4.72 $\pm$ 0.79 4.03 $\pm$ 0.26 4.208$\pm$ 0.061 5.119 $\pm$ 0.162
Lens Mass ($M_{\odot}$) 0.11 $\pm$ 0.01 0.12 $\pm$ 0.02 $0.64^{+0.22}_{-0.34}$ 0.661 $\pm$ 0.053 0.235 $\pm$ 0.023
Lens Distance (kpc) 0.81 $\pm$ 0.10 1.24 $\pm$ 0.18 $4.6^{+1.1}_{-1.8}$ 4.412 $\pm$ 0.262 1.813 $\pm$ 0.205
Planet Mass ($M_{\rm Earth}$) 9.2 $\pm$ 2.2 15.2 $\pm$ 5.9 $109^{+38}_{-58}$ 169.5 $\pm$ 46.04 24.75 $\pm$ 3.39
$\alpha_{\perp}$(AU) 0.92 $\pm$ 0.16 1.21 $\pm$ 0.27 $3.8^{+0.9}_{-1.5}$ 3.775 $\pm$ 0.204 1.990 $\pm$ 0.164

### v2

@@ 4 Image Constrained Light-Curve Model
In Table 4, we show the light-curve model parameters with the values and uncertainties for the Einstein radius crossing time ($t_{\rm E}$), time of closest approach ($t_{0}$), source crossing time ($t_{*}$), impact parameter ($u_{0}$), the instantaneous projected separation which is scaled to $\theta_{\rm E}$ ($s$), the host-planet mass ratio ($q$), and the angle between the source trajectory and traverse line that passes through the two lenses ($\alpha$). The model parameters that are essential for obtaining solutions to the lens system’s physical parameters are $q$, $_{\rm E}, tt_{*}$, and the microlensing parallax $\pi_{\rm E}$.
Another essential parameter is the source distance $D_{\rm S}$, which is needed to determine the mass of the lens from the relative proper-motion constraint, as shown in equation 5. Therefore, we include $D_{\rm S}$ as a light-curve model parameter and use the Koshimoto et al. (2021) Galactic model as a prior.
The Galactic bulge is located close to the ecliptic plane therefore microlensing events with significant microlensing parallax signals often exhibit an “ecliptic degeneracy” (Poindexter et al., 2005) which involves replacing the microlensing geometry with its mirror image. In the case of this event, the directions of relative proper motion, $\mu_{\rm rel}$, differs by more than $160^{\circ}$, so the Keck $\mu_{\rm rel}$ measurement removes this degeneracy.
[…]
To investigate this degeneracy and understand how the light-curve model parameters are being affected, we make a diagonal cut in the ($\pi_{\rm E,E},\pi_{\rm E,N}$) space on the corner plot (Figure 7), at $\pi_{\rm E,E}=\pi_{\rm E,N}-0.33$. In Table 4 we present the light-curve model parameters associated with the “large” and “small” parallax values, where “large” refers to $\pi_{\rm E,N}>\pi_{\rm E,E}-0.33$, and “small” refers to $\pi_{\rm E,N}<\pi_{\rm E,E}-0.33$. We also do this in Table 5 while presenting the lens system’s physical parameters. In Section 5.1 we discuss this degeneracy further.
4.1 Source Color and CMD
Using the measured total magnitude and color from HST as a constraint on the light-curve model, we find MCMC distributions of the primary source’s (S1) magnitude and color to be:
$(V-I,I)_{\rm S1}=(2.188,19.613)\pm(0.038,0.021)\ \ \ \ {\rm for\ small}\ \pi_{\rm E},$
$(V-I,I)_{\rm S1}=(2.190,19.617)\pm(0.030,0.017)\ \ \ \ {\rm for\ large}\ \pi_{\rm E},$
and for the secondary source (S2):
$(V-I,I)_{\rm S2}=(2.036,20.807)\pm(0.502,0.244)\ \ \ \ {\rm for\ small}\ \pi_{\rm E},$
$(V-I,I)_{\rm S2}=(1.894,20.846)\pm(0.324,0.183)\ \ \ \ {\rm for\ large}\ \pi_{\rm E}.$
To calculate the source angular size of the primary source, we need the dereddened magnitude and color. The extinction to the source star is determined by using the red clumb stars within 120” of the MB10328 event from the OGLE-III catalogue (Szymański et al., 2011). Therefore, we find $A_{I}=1.817$ and $A_{V}=3.247$. Using the following relation from Boyajian et al. (2014):
$\log_{10}[2\theta_{*}/(1\rm mas)]=0.5014+0.4197(V-I)_{\rm s0}-0.2I_{\rm s0},$ (3)
we calculate the primary source’s angular radius $\theta_{*}=0.910\pm 0.042\ \mu\rm as$ (for small $\pi_{\rm E}$) and $\theta_{*}=0.911\pm 0.032\ \mu\rm as$ (for large $\pi_{\rm E}$ ). We can estimate the relative source-proper motion using $\mu_{\rm rel,G}=\theta_{*}/t_{*}=3.587\pm 0.436$ mas $\rm yr^{-1}$ (for small $\pi_{\rm E}$) and $\mu_{\rm rel,G}=5.099\pm 0.243$ mas $\rm yr^{-1}$ (for large $\pi_{\rm E}$) , where the notation $G$ represents the geocentric reference frame. This cannot be directly compared to the relative proper motion measured in Keck, since Keck is in the heliocentric reference frame. In Table 5 we show our converted Keck proper motion, after converting it to the geocentric reference frame (see Section 5 for details).
From the light-curve model we also find the MCMC distributions of the lens $V$ and $I$ measurements:
$(V-I,I)_{\rm L}=(3.036,21.551)\pm(0.743,0.412)\ \ \ \ {\rm for\ small}\ \pi_{\rm E},$
$(V-I,I)_{\rm L}=(3.756,22.318)\pm(0.193,0.133)\ \ \ \ {\rm for\ large}\ \pi_{\rm E}.$
It should be noted that these are estimates from the light-curve model, which includes a constraint on the total flux on the source and lens from the HST data. The lens color and magnitude from the light curve model are not significantly different from the values measured for the lens in the HST follow-up data (see Table 3). We plot these values, and of the two sources, on a color-magnitude diagram (CMD) in Figure 8. The degeneracy in the microlensing parallax does not have a significant effect on the primary source’s color and magnitude. However, the secondary source and the lens are impacted by the degenaracy, as seen on Figure 8.
The higher order effects that were included in the light-curve modeling of this event are discussed further in Section 4.2.
4.2 Higher-order effects
@@ 4.2.2 Orbital Motion of Lens System
Orbital motion of the lens system is always present, however, their positions are often assumed to be fixed during the microlensing event. This is because a microlensing event’s duration is often on the order of a few days to weeks, making is a small fraction of the lens systems’ orbital period of several years. In our model we assume a face-on circular orbit with the parameters of orbital angular speed, $\omega$, and the binary separation rate of $ds/dt$. If microlensing parallax effects are present on the light-curve, then it is physically realistic that the model include orbital motion effects too. However, due to rotational symmetry, $\pi_{\rm E,E}$ (which is the component perpendicular to the instantaneous direction of the Earth’s acceleration) is often degenerate with $\omega$ (Batista et al., 2011; Skowron et al., 2011).
@@ 4.2.3 Binary Source
The parameters due to the binary source effect are the time of closest approach ($t_{0,2}$) and impact parameter ($u_{0,2}$) of the second source. Additionally, we have the $I$ and $V$ band flux ratios between the second source ($S_{2}$) and the primary source ($S_{1}$), fr2I and fr2V, respectively. The xallarap effect is when the source orbital motion due to the second source invokes the parameters $\rm dt_{\rm E21}$, which is implied by $t_{\rm E2}-t_{\rm E1}$; and $d\alpha$, implied by $\alpha_{\rm E2}-\alpha_{\rm E1}$. In the light-curve modeling we also include the inverse of the source star orbital period, $1/T_{\rm Sbin}$.
Table 4: Light-curve model Parameters using the microlensing parallax plus orbital motion close and wide solution, and the xallarap solution from Furusawa et al. (2013); and light-curve model parameters from this work, separated into the ‘small’ and ‘large’ $\pi_{\rm E}$ parallax degeneracy. The new model includes microlensing parallax, orbital motion, xallarap, and magnification of a second source. Parameter descriptions can be found throughout Section 4.
Parallax plus orbital motion This Work
Parameter ($u_{0}<0$) ($u_{0}>0$) Xallarap small $\pi_{\rm E}$ large $\pi_{\rm E}$
$t_{\rm E}$ (days) $62.6\pm 0.6$ $64.2\pm 0.6$ $61.8\pm 0.3$ $61.24\pm 1.37$ 63.51 $\pm$ 0.84
[…]
$u_{0}$ $0721\pm 0.0008 0.0$$716\pm 0.0007 0.0$$741\pm 0.0004 0.0$$661\pm 0.0019 0.0$$0507\pm 0.0019$
s $154\pm 0.016 1.1$$180\pm 0.028$ $1.220\pm 0.012$ 1.222 $\pm$ 0.006 1.228 $\pm$ 0.004
[…]
$\alpha$ $0.2743\pm 0.0087$ 0.1965 $\pm$ 0.0151 $0.2024\pm 0.0004$ $2.9937\pm 0.0011$ $3.0272\pm 0.0047$
$\pi_{\rm E,N}$ 1.01 $\pm$ 0.06 0.72 $\pm$ 0.05 0.138 $\pm$ 0.019 0.400 $\pm$ 0.036
$\pi_{\rm E,E}$ $0.51\pm 0.04$ $39\pm 0.03 0.0$$007\pm 0.011$ $0.247\pm 0.056$
$ds/dt\times 10^{3}\ (\rm days^{-1})$ $2.51\pm 0.63$ $1.41\pm 1.16$ $0464\pm 0.0864 0.0$$0915\pm 0.0288$
$\omega\times 10^{3}\ \rm(rad/day)$ $7.39\pm 0.39$ $1.39\pm 0.60$ $0.25\pm 1.05$ $2.60\pm 0.30$
[…]
$u_{0,2}$ $0.210\pm 0.989$ 1.201 $\pm$ 0.099
[…]
$\rm d\alpha$ (mas) $0115\pm 0.0584 0.0$$0126\pm 0.0368$
[…]
$t_{\rm fix}$ (days) - - - 5406.0 5406.0
$D_{\rm S}$ (kpc) 7.680 $\pm$ 0.575 7.786 $\pm$ 0.516
$\chi^{2}$ 5657.75 5660.31 5652.59 4654.90 4654.28
dofdof 5660 5660 5658 4682 4682
Figure 7: Corner plot from the MCMC probability distribution showing the parameters due to the microlenisng parallax, orbital motion, xallarap and magnification of the secondary source.
Figure 8: CMD of OGLE III stars within 120 arseconds of the MB10328 microlensing event, in grey. In orange is the color-magnitude diagram of the HST field from the 2018 epoch transformed to the same extinction and Galactic bar distance. The red clump giant centroid is indicated with a red dot. The purple marker is the primary source (S1), and the blue markers show the secondary source (S2) color and magntitude. The green markers represent the lens. The diamond markers indicate the values for the small $\pi_{\rm E}$ distribution, while the circles indicate values for the large $\pi_{\rm E}$. For S1, the difference and the error bars are too small to be seen on the figure.
5 Planetary System Parameters
Once the best-fit model parameters have been found, we can sum over the MCMC results presented in Table 4, and determine the posterior distribution of the physical properties of the lens system which are presented in Table 5. We use light-curve parameters, in combination with the mass-distance and mass-luminosity relations in Equations 4, 5, and 8. The constraints from the high-resolution follow-up data have already been applied in the light-curve modeling, and are therefore not applied again while summing over the MCMC results. Additionally, the source distance, $D_{\rm S}$, prior is also applied during the light-curve modeling, and therefore, not applied again. However, a Galactic model prior is applied on the lens distance for the $D_{\rm S}$ value.
A common assumption used in the Bayesian analysis of microlensing events is that all stars are equally likely to host a planet with a fixed mass ratio, $q$. However, more recent studies (Koshimoto et al. (2021); Bennett et al. (2024)) show that planet hosting probability of a fixed mass ratio increases linearly with the host mass. Therefore, this prior is included when summing over the MCMC results. Since we have constraints from high resolution imaging, this prior has a small effect on the results.
@@ 5 Planetary System Parameters
Two mass-distance relations can be used which include the angular Einstein ring radius, $\theta_{\rm E}$, or the microlensing parallax $\pi_{\rm E}$ (Beaulieu et al., 2008; Gaudi, 2012). These are:
@@
where $\pi_{\rm rel}=\rm AU(\frac{1}{D_{\rm L}}-\frac{1}{D_{\rm S}}),$ and $\kappa=4G/(c^{2}AU)=8.144\ \rm mas\ M_{\odot}^{-1}$. Therefore, with Equations 4 and 5 an expression can be found for the lens mass $M_{\rm L}$, with no dependance on the lens or source distance, $D_{\rm L}$ or $D_{\rm S}$, respectively. The lens distance, however, does depend on the source distance.
For the event MB10328, a direct measurement of the lens-source proper motion has been made, therefore $\theta_{E}$ can be constrained with $\theta_{E}=\mu_{\rm rel,G}t_{E}$. However, $\mu_{\rm rel,G}$ is in the instataneous geocentric inertial reference frame which moves with the earth at peak magnification. Since the proper motion derived from the high angular resolution follow up images are in the heliocentric frame, it needs to be converted to the geocentric one. This can be done using the equation:
${\mu}_{\rm rel,G}=\mu_{\rm rel,Hel}-\frac{v_{\oplus}\pi_{\rm rel}}{\rm AU}$ (6)
Where $v_{\oplus}$ is the projected velocity of the earth relative to the sun at the time of the events peak magnification. For MB10328, this was $v_{\oplus E,N}=(28.878,-0.689)$ km/sec =(6.088,0.145)=(6.088,-0.145) AU/yr at HJD’ = 5406. Since $\pi_{\rm rel}$ is the relative parallax, we can rewrite equation 6 as:
${\mu}_{\rm rel,G}=\mu_{\rm rel,Hel}-(6.088,-0.145)\times(1/D_{\rm L}-1/D_{\rm S}),$ (7)
where $\mu_{\rm rel,Hel}$ and $\mu_{\rm rel,G}$ are in units of mas/yr and $D_{\rm L}$ and $D_{\rm S}$ are in units of kpc. This relation is used in our Bayesian analysis of the light-curve (Section 4). Therefore, from our Keck data we determine the geocentric relative proper motion to be $\mu_{\rm rel,G}=4.208\pm 0.061\ \rm mas\ yr^{-1}$ (for small $\pi_{\rm E}$), and $\mu_{\rm rel,G}=5.119\pm 0.162\ \rm mas\ yr^{-1}$ (for large $\pi_{\rm E}$).
The angular Einstein radius, $\theta_{\rm E}$, can be calculated by two methods. First by using light-curve model parameters, $\theta_{\rm E}$ can be calculated with $\theta_{\rm E}=\theta_{*}/\rho$, where $\rho=t_{*}/t_{\rm E}$ and $\theta_{*}$ are found from the color-surface brightness relation in Equation 3; and secondly, from the measured lens-source proper motion in the follow-up data, once it has been converted to the geocentric reference frame.The final $\theta_{\rm E}$ measurement, presented in Table 5, comes from the weighted sum of these two methods, using the lens-source proper motion measured from Keck; the lens-source relative proper motion light-curve model parameters $t_{*}$, $t_{\rm E}$; and the source $V$ and $I$ magnitude using a color-surface brightness relation (Boyajian et al., 2013).
@@
$m_{\rm L}=10+5\log_{10}(D_{\rm L}/1\ \rm kpc)+A_{K,L}+M_{\mathcal{K}}(M_{L}),$ (8)
[…]
By summing over the image-constrained light-curve model parameters and using the mass-distance and mass-luminosity relations, we determine the posterior distribution of the physical properties of the lens system, which are presented in Table 5. However, as we expected the distribution of the lens system’s physical parameters is not a single Gaussian, and when looking at the probability distributions in Figure 9 we see a clear degeneracy, which we investigate in more detail in the following section 5.1.
Table 5: Planetary system parameters from high angular resolution follow up constraints compared to those found by Furusawa et al. (2013).We show the lens system parameters for the large and small $\pi_{\rm E}$ values.
. Parallax plus orbital motion This Work Parameter ($u_{0}<0$) ($u_{0}>0$) Xallarap small $\pi_{\rm E}$ large $\pi_{\rm E}$ $\theta_{\rm E}$ (mas) 0.98 $\pm$ 0.12 0.83 $\pm$ 0.14 0.68 $\pm$ 0.04 0.713 $\pm$ 0.021 0.881 $\pm$ 0.347 $\mu_{\rm rel,G}$ ($\rm mas\ yr^{-1}$) 5.71 $\pm$ 0.70 4.72 $\pm$ 0.79 4.03 $\pm$ $246±\pm 0.$068 5.067 $\pm$ 0.191 Lens Mass ($M_{\odot}$) 0.11 $\pm$ 0.01 0.12 $\pm$ 0.02 $0.64^{+0.22}_{-0.34}$ 0.633 $\pm$ 0.062 0.244 $\pm$ 0.030 Lens Distance (kpc) 0.81 $\pm$ 0.10 1.24 $\pm$ 0.18 $4.6^{+1.1}_{-1.8}$ 4.350 $\pm$ 0.320 1.928 $\pm$ 0.269 Planet Mass ($M_{\rm Earth}$) 9.2 $\pm$ 2.2 15.2 $\pm$ 5.9 $109^{+38}_{-58}$ 141.13 $\pm$ 42.42 24.75 $\pm$ 3.39 $\alpha_{\perp}$(AU) 0.92 $\pm$ 0.16 1.21 $\pm$ 0.27 $3.8^{+0.9}_{-1.5}$ 3.787 $\pm$ 0.242 2.075 $\pm$ 0.210
Figure 9: Bayesian probability distributions of the planetary system, including host mass and distance, separation and planetary companion mass. Orange represents $\pi_{\rm E,N}>\pi_{\rm E,E}-0.33$ values, and green respresents $\pi_{\rm E,N}<\pi_{\rm E,E}-0.33$ values. The central 68.3% of the distributions are shaded in darker hues (of either orange or green) and the remaining central 95.4% of the distributions are shaded in lighter hues. The vertical black line marks the median of the probability distribution of the respective parameters for the large and small $\pi_{\rm E}$ values.